OpenStudy (anonymous):

The length of a rectangle is 2 inches more than twice its width, and its perimeter is 60 inches. Find the length and width of the rectangle. If w = the width of the rectangle, then the length equals _____.

OpenStudy (anonymous):

Well you can setup some equations... So: Length=2+2*width Width=width Perimeter=60=2+2w+2+2w+w+w (This is adding the two length and two width sides) That should get you on the right track, just solve for w or the width and go from there :)